The vapor pressure of pure water at 26°C is 25.2 torr. What is the vapor pressure of a solution which contains 20.0 g of glucose (C6H12O6), in 70 g water?


1
Expert's answer
2021-11-18T09:31:02-0500

Number of moles of glucose =20/180=0.111

Number of moles of water =70/18= 3.89

Mole fraction of glucose= 0.111/0.111+3.89=0.0278


The relative lowering in vapour pressure is equal to the mole fraction of solute.

25.21-p/25.21= 0.0278

P=25.21−(25.21×0.0278)=24.5 torr